Let A be a square matrices of order
m = min no of zero if A is a triangular matrices
p = min no of zero if A is a diagonal matrices
l = max no of different entries if A is a upper triangular matrix

Grade:12

1 Answers

sudhir pal
10 years ago
your question is not very clear so i'm taking some assumption
m = order of matrix
n =min no of zero if A is a triangular matrices = m(m-1)/2
p =min no of zero if A is a diagonal matrices= m(m-1)
l = max no of different entries if A is a upper triangular matrix = m(m+1)/2
I hope you got your answer
Thanks & Regards
Sudhir,
askIITians Faculty
Qualification.
IIT Delhi